# Collaborative Learning in Agentic Systems: A Collective AI is Greater Than the Sum of Its Parts

### v1

@@ 2.3 Knowledge Selection and Sharing
Phase 1: Query-Metadata.
At any time, an agent $i$ can initiate a task embedding query (TEQ) by broadcasting its current task embedding $v_{i}$, iteration performance $r_{i}$, and its IP/port (Figure 1(B)). Upon receiving the TEQ, each peer agent $j\neq i$ responds with its most recent task embedding $v_{j}$ and associated performance score $r_{j}$.
Phase 2: Mask Selection and Acquisition.
Once the agent has received query responses (QR) with relevant embeddings from other agents, it computes the cosine similarity between its own embedding and those of each peer,
@@
In practice, cosine similarities between Wasserstein task embeddings are typically positive. Cosine similarity was observed to perform better than Euclidean distance in preliminary tests, and showed scale-invariance and empirical robustness in high-dimensional settings. To identify useful policies, agent $i$ applies two heuristic filters based on similarity and performance:
$\mathbb{I}_{\text{align}}(i,j)$ =$=\begin{cases}1&\text{if }d_{\text{cos}}(v_{i},v_{j})>\theta,\\ 0&\text{otherwise}\end{cases}$ (Criterion 1)
$\mathbb{I}_{\text{perf}}(i,j)$ =$=\begin{cases}1&\text{if }r_{j}>r_{i},\\ 0&\text{otherwise}\end{cases}$ (Criterion 2)
with $\theta=$ set to $0.5$ in all experiments. The first criterion promotes semantic alignment between tasks, revealing latent similarities that may lead to effective transfer (see Figure 4). The second ensures that agents only acquire masks from peers whose performance exceeds their own, avoiding noisy or under-trained policies that can degrade performance. Criterion 2 guards against the frequent false positives seen when comparing poorly trained policies that yield deceptively similar embeddings. The selection criteria can also be used on the peers after they receive the query. This method acts as a bandwidth-efficient alternative that enables communication without affecting the overall outcome of the approach. Additionally, this method offers a scalable pathway for extensions to lifelong learning, which are discussed in Appendix B.
The peer masks that pass both criteria are stored as the set $P_{c+1}=\{\phi_{1},...,\phi_{K}\}$, where each $\phi_{k}$ denotes a mask received from a selected peer. This two-phase strategy ensures that agents acquire knowledge that is not only relevant but also demonstrably effective.
2.4 Knowledge Composition
Masks received from other agents meeting the two selection criteria above have been pretrained on tasks that are likely related to, or share similarities with, the agent’s current task. As originally shown in Ben-Iwhiwhu et al. (2023), a linear combination of masks, weighted with trainable parameters $\beta$, can lead to beneficial policy search in RL. MOSAIC exploits this idea by combining policies that have been trained on the local agent, plus those received from other agents. In this study, the $\beta$ values are computed via softmax over a set of beta parameters optimized in log-space, $\tilde{\beta}\in\mathbb{R}^{1+|P_{c}|}$, where $P_{c}$ is the number of masks received. During the forward pass, the agent constructs the linearly combined, binarized mask,
@@ 2.4 Knowledge Composition
where $\phi_{\tau}$ is the agent’s own task mask, and $\phi_{k}$ is the $k$-th peer mask. The parameters $\beta_{\tau}$ and $\beta_{k}$ are the softmax-normalized values derived from $\tilde{\beta}$, for $\phi_{\tau}$ and $\phi_{k}$, respectively. The resulting binary mask, $\phi_{\tau}^{\text{lc}}$, is then used to modulate the backbone parameters, yielding the final policy $\pi_{\tau}$. Using the STE method, the agent updates the $\beta$ parameters and $\phi_{\tau}$ via backpropagation, keeping all the $\phi_{k}$ masks fixed.
Reward-Guided Initialization.
The trainable $\beta$ parameters are ideal to enable gradient descent to discover which policy is most beneficial on the current task. However, given the available information, one question is: What is a good initial weighting for the set of masks that include the local mask and those received from other agents? MOSAIC proposes a reward-guided weighting mechanism to set the initial values. While such values can be later tuned by gradient descent, an ablation study suggested that performing a good initial guess leads to a significant advantage. The intuition is that an agent with low performance may favor external knowledge, while an agent with higher returns could rely more on their own policies. The initial set of $\beta$ is given by
@@
where $r\in[0,1]$ is the agent’s normalized average return over the most recent iteration and $|P_{c+1}|$ is the number of peer masks received during a communication event. This scheme encourages a smooth shift from exploration to self-reliance as learning progresses.
The process just described occurs multiple times, and specifically, each time new masks are received. Therefore, before a new set of peer masks $P_{c+1}$ is integrated, the agent performs a consolidation operation that combines its current task mask $\phi_{\tau}$ with the set of previous peer masks $P_{c}$. This is done using a weighted linear combination similar to the forward pass. The resulting consolidated mask updates the existing $\phi_{\tau}$,

### v2

@@ 3.3 Knowledge Selection and Sharing
Phase 1: Embedding queries
At any time, an agent $i$ can initiate a task embedding query (TEQ) by broadcasting its current task embedding $v_{i}$, iteration performance $r_{i}$, and its IP/port (see Figure 1(B)(C)). Upon receiving the TEQ, each peer agent $j\neq i$ responds with its most recent task embedding $v_{j}$ and associated performance score $r_{j}$.
Phase 2: Policy selection.
Once the agent has received query responses (QR) (see Figure 1(F)(G)) with relevant embeddings from other agents, it computes the cosine similarity between its own embedding and those of each peer,
@@
To identify useful policies, agent $i$ applies two heuristic filters based on similarity and performance:
$\mathbb{I}_{\text{align}}(i,j)$ $=\begin{cases}1&\text{if }d_{\text{cos}}(v_{i},v_{j})>\theta,\\ 0&\text{otherwise}\end{cases}$ (Criterion 1)
$\mathbb{I}_{\text{perf}}(i,j)$ $=\begin{cases}1&\text{if }r_{j}>r_{i},\\ 0&\text{otherwise}\end{cases}$ (Criterion 2)
with $\theta=0.5$ in all experiments. The first criterion promotes semantic alignment between tasks, revealing latent similarities that may lead to effective transfer (see Section 4.4). The second ensures that agents only acquire masks from peers whose performance exceeds their own, avoiding noisy or under-trained policies that can degrade performance. Criterion 2 guards against the frequent false positives seen when comparing poorly trained policies that yield deceptively similar embeddings.
The two phases and the selection criteria also result in a bandwidth-efficient approach as policies are transferred only when considered potentially useful. The peer masks that pass both criteria are stored as the set $P_{c+1}=\{\phi_{1},...,\phi_{K}\}$, where each $\phi_{k}$ denotes a mask received from a selected peer.
3.4 Knowledge Composition and Fine Tuning
Masks received from other agents meeting the two selection criteria above have been pretrained on tasks that are likely related to, or share similarities with, the agent’s current task. As previously shown in Ben-Iwhiwhu et al. (2023), a linear combination of masks, weighted with trainable parameters $\beta$, can lead to beneficial policy search in RL. MOSAIC exploits this idea by combining policies that have been trained on the local agent, plus those received from other agents. In this study, the $\beta$ values are computed via softmax over a set of beta parameters optimized in log-space, $\tilde{\beta}\in\mathbb{R}^{1+|P_{c}|}$, where $P_{c}$ is the number of masks received. During the forward pass, the agent constructs the linearly combined, binarized mask,
@@ 3.4 Knowledge Composition and Fine Tuning
where $\phi_{\tau}$ is the agent’s own task mask, and $\phi_{k}$ is the $k$-th peer mask. The parameters $\beta_{\tau}$ and $\beta_{k}$ are the softmax-normalized values derived from $\tilde{\beta}$, for $\phi_{\tau}$ and $\phi_{k}$, respectively. The resulting binary mask, $\phi_{\tau}^{\text{lc}}$, is then used to modulate the backbone parameters, yielding the final policy $\pi_{\tau}$. As training continues on the local task, the agent updates the $\beta$ parameters and $\phi_{\tau}$ via backpropagation (keeping all the $\phi_{k}$ masks fixed) thus fine tuning the overall resulting policy determined by $\phi_{\tau}^{\text{lc}}$.
Reward-Guided Initialization (RGI).
The trainable $\beta$ parameters enable gradient descent to determine which policies are most useful for the current task. The initial weighting as the masks are received is given by
@@
where $r\in[0,1]$ is the agent’s normalized return from the last iteration and $|P_{c+1}|$ is the number of received peer masks. This scheme biases low-performing agents toward external knowledge and high-performing agents toward their own policies. Ablation studies show that such an initialization provides a strong advantage, even though $\beta$ is later tuned by gradient descent.
Before integrating a new set of peer masks $P_{c+1}$, the agent consolidates its current task mask $\phi_{\tau}$ with the previous peer masks $P_{c}$ using a weighted linear combination,